OpenStudy (anonymous):

The quotient of a number and five is eight less than the number. What is the number?

OpenStudy (anonymous):

call the number n n/5 = n-8 n = 5n - 40 40 = 4n n = 34982395783457398239858738945

OpenStudy (anonymous):

thank you @JamesWolf